# The Conspiracy Money Machine: Uncovering

### v2

@@ 4.3.1 Longitudinal analysis
An interesting aspect to explore is how conspiracy communities have evolved over time. We analyze this dimension in Fig. 2, which shows the number of channels created daily on Telegram. The Figure is divided into five parts: the first four charts ($a$,$b$,$c$,$d$) analyze each community, while the last one ($e$) compares the four conspiracy communities aggregated against the other Telegram communities. The chart shows that the creation of conspiracy channels is not evenly distributed over time. Instead, we find two spikes in channel creation. The first one begins around mid-March 2020, reaches a peak in May, and starts declining until June 2020. Instead, the second spike is much steeper and goes from 2021-01-06 to the middle of March 2021.
@@ 5.1 Donation platforms
Total 123,129 2,819 1,175,557 136,710
The first monetization strategy we analyze consists of asking the subscribers for donations to support the channel and its activity. As a first step, we identify the most popular donation platforms that can be used for this purpose. We manually collect them by analyzing the results of Google queries such as: top donation platforms. Given the presence of language specific communities, we conducted country-specific queries containing different languages and keywords. Moreover, to avoid search results on Google from being affected by our browsing history or geo-location, we conduct queries using a VPN. In this way, we collect 31 donation services. Then, we analyze the presence of donation platforms in the URLs shared in the Extended Conspiracy URLs Dataset.
@@
After filtering out messages classified by the LLM as discrediting the linked donation page, we identified 124,610 URLs from 15 services, shared across 5,585 channels, accounting for 31.3% of our dataset. The most used platforms by number of URLs are Paypal/donate [83] with 83,512 URLs (66% of the donation URLs), Patreon with 31,041 URLs (24.5% of the donation URLs), BuyMeACoffee with 3,157 (2.5%), SubScribeStar with 2,604 (2.1%), DonorBox with 1,550 (1.2%) and Ko-Fi with 1,475 (1.2%).
To estimate the potential earnings of the Conspiracy Channel Dataset on these platforms, we create a custom crawler using the Selenium Framework [90]. We crawl all the URLs from the 6 most used donation platforms in our dataset, covering 98.9% of the URLs.  [^1] We conducted the crawling of donation, crowdfunding, and affiliate URLs in September 2023. Unfortunately, PayPal/donate and DonorBox do not provide information about donation amounts and the number of donors. Thus, by scraping the web pages of these services, we could only determine the number of unique profiles still reachable. Instead, we obtained information regarding the number of donors and donated amounts for most profiles on the other services. Specifically, for Patreon and SubscribeStar, we retrieve the number of donors and the total amount of money donated in the last month. As for BuyMeACoffee and Ko-Fi, we collect data on the number of donors and the total amount of money raised by the campaigns’ creators.
@@ 5.2 Crowdfunding and Fundraising services
Thus, we use the same approach as the previous subsection to discover these cases (Sec. 5.1), asking the GPT-4o model to determine the author’s intention behind each message (prompt in Appendix, Tab. 8). Out of 7,499 distinct messages containing a URL to a crowdfunding project, the model classifies 7,319 messages (97.6%) as promoting campaigns while 180 as messages against the linked campaign. Also in this case, to validate the model, we randomly sampled 200 messages, 100 marked as positive and 100 as negative. Then, we manually label these messages as previously done in Sec. 5.1. Here, we identified 145 positive samples and 55 negative samples. Comparing the model results with the manually annotated labels, we observed the model tends to classify positive messages as negative, similar to the previous task. Specifically, the model incorrectly classified 45 messages endorsing crowdfunding campaigns as negative. After reviewing the misclassified messages, we find many authors initially discredit the GoFundMe platform, encouraging their audience not to use it, and then endorse a crowdfunding campaign on the GiveSendGo platform. This pattern was common among American far-right channels, which frequently complain about GoFundMe blocking their projects. Given that the model provides a conservative estimation and considering the relatively low number of misclassified messages (180), we used this model for our classification. Accordingly, we discarded 180 unique messages (357 including duplicates) linking to crowdfunding platforms.
As further validation, we manually examine the crowdfunding campaigns that raised more than \$200K, counting 98 different projects, to determine if they are related to conspiracy theories. Three researchers independently analyzed these projects, and a campaign was labeled as related to conspiracy theories only if there was unanimous agreement. Among the 98 campaigns, we identified 51 projects (52.04%) that, despite being supported by conspiracy channels, have goals unrelated to conspiracy theories.
@@
Tab. 4 reports the number of URLs, different projects, money raised, and number of backers for the top eight services by number of URLs. As it is possible to note, this strategy is the most remunerative, collectively funding conspiracy theorists with almost \$66M. Analyzing the content of the campaigns and find that they typically fall into these categories:
Campaigns supporting far-right. We identified several fundraising campaigns for far-right projects, mainly raising funds for legal costs related to the January 6th Capitol riot. Three of these campaigns alone  [58, 87, 57] collected \$160,816 from 2,891 backers. We find that many of these campaigns are hosted on Givesendgo, a platform well known for promoting extremist content [6].
Campaigns about COVID-19. We also find GiveSendGo campaigns frequently linked to COVID-19, including those for the Freedom Convoy 2022 [51], a protest movement against COVID-19 policies. Similar campaigns on GoFundMe redirect to refund pages due to the platform’s ban on content promoting violence [50]. Looking at the news, we discover that one of these campaigns was the biggest gainer of the platform, raising over \$10 million from 120,000 donors [19]. Additionally, Indiegogo and Kickstarter host projects funding documentaries about COVID-19 [32, 94, 105]. These platforms have stricter content policies, but moderating pseudo-scientific claims remains challenging due to freedom of expression concerns.
Scam campaigns. We detect crowdfunding campaigns that are outright scams. An example is a Fundrazor campaign [69] falsely claiming to raise funds for starving children in Venezuela. It collected \$39,500 from 602 donors before being shut down, and a banner on the page states that Save The Children has confirmed they have no association with it. Another scam [5] on Indiegogo raised over \$7.4K for a portable air cleaner claiming to purify the air with a high-frequency generator before being closed as the creator claimed that Silicon Valley internet companies had hindered the project.
@@ 5.3 Affiliate Programs
Conspiracy channels adopting this strategy use affiliate programs from e-commerce platforms to earn a commission. Thus, as a first step, we identify the most popular e-commerce platforms they can utilize for this purpose. To this end, we leverage SimilarWeb to retrieve all the 42 services reported in the global rank for the eCommerce & Shopping category. Then, we extract URLs containing the domain of these platforms from the Extended Conspiracy URLs Dataset, finding 70,149 URLs from 28 platform. Among the top 4 platforms, covering 98.1% of e-commerce links, only Amazon and eBay offer an affiliate program. In this program, a participant, known as partner, can generate unique links pointing to products on the platform that embed his unique identification number. The partner earns a commission for any purchases made by users who land on the platform through his link. In particular, Amazon’s partners earn a commission, between 1% and 12% depending on the categories of the product and the location of the targeted market (e.g., amazon.com, amazon.de) from all the items bought in the next 24 hours from arrival through the link [10]. Similarly, on eBay, a partner earns a commission between 1% and 4% on the items purchased on the platform in the following 24 hours [43].
Overall, we find 63,445 URLs pointing to Amazon and 1,932 to eBay in the Conspiracy Channel Dataset. To detect URLs belonging to the affiliate program, we search for URLs containing the parameters tag= for Amazon and campid= for eBay. Surprisingly, we discover that 34,412 (54.2%) of Amazon’s URLs and 131 (6.8%) on eBay are affiliated links granting commissions to the conspiracy channels and that 1,810 channels (10.2% of our dataset) leverage this strategy. Tab. 5 provides a summary of the metrics of e-commerce platforms.
@@ 5.4 Monetization Analysis
Monetization links promoted by multiple channels. The analysis of Sabmyk affiliate links highlights the sharing of the same referral link by multiple channels. To study this phenomenon in more depth, we conducted a broader analysis that also includes donations and crowdfunding projects. Fig. 3 shows the cumulative distribution function (CDF) depicting the fraction of referral IDs (orange), crowdfunding (blue), and donation projects (green) shared in channels. The main plot provides an overview of the entire distribution, while two subplots show two different levels of zoom. By examining the zoomed CDFs, we can observe that most referral links, donation projects, and crowdfunding projects are typically shared by only a few channels. Specifically, we find that 65.4% of affiliate IDs, 56.8% of donation projects, and 55.8% of crowdfunding projects are shared by only a single channel. The figure also highlights that affiliate links are shared by fewer channels than donation and crowdfunding projects, with almost 80% of referral links shared by at most two channels. This can be expected as, while affiliate links are primarily beneficial to the channel owner, crowdfunding and donation projects may sometimes support broader causes and may be shared by channel owners even if they are not the direct beneficiaries of the funds received. Looking at the tail of the overall distribution, we find other interesting insights. Examining crowdfunding projects, we find that the most widely shared is a highly popular project supporting the Freedom Convoy, promoted on GiveSendGo [104]. This project is shared by 247 conspiracy channels and has raised \$9.7 million. For donation projects, we identified a donation link that was shared by 844 channels. This link is associated with the Corona_Fakten channel, which focuses on disseminating news and controversial theories about COVID-19. The channel frequently includes the donation link in its posts, and because its content is widely shared and forwarded by many other channels, the donation link has spread across hundreds of channels. What is more surprising instead, is the presence of a group of over 549 channels promoting Amazon products using the same affiliate ID. Our analysis reveals this group of channels is entirely located in the German community and is actively involved in promoting various conspiracy theories, predominantly related to COVID-19.
@@
Shift in platform usage. Fig. 7 and Fig. 8, in the Appendix, show, respectively, the proportion of donation and crowdfunding platforms used by each community monthly. Examining donation platforms, we observe distinct preferences across different linguistic communities. The German community predominantly uses PayPal. In contrast, Patreon has been the most used platform in the Neo-Latin community over the last year. The English community initially used Patreon and PayPal most frequently, but in the past 18 months, there has been a significant diversification, with multiple platforms being utilized. For crowdfunding, PayPal Pools was the dominant platform among both the German and Neo-Latin communities. However, its popularity decreased until it was discontinued in September 2021, replaced by GoFundMe. The English community displayed a different trend. GoFundMe was consistently the preferred platform, as PayPal Pools was never widely adopted. Around October 2020, GiveSendGo began to rise in popularity, accounting for 75% to 90% of all URLs shared by the English community monthly. This shift may be attributed to GoFundMe stricter policies, as GiveSendGo promotes itself as having no moderation. We find evidence of this transition in the messages of the English community.
@@ 5.5 Other Strategies
In addition to the previously analyzed monetization strategies, we find conspiracy channels can exploit other sources of revenue.
@@
Blockchain addresses. In our analysis, we discover conspiracy channels asking for cryptocurrency donations in their descriptions. We use regular expressions to extract wallet addresses of the most popular blockchains (Bitcoin and Ethereum), as well as the prominent privacy-preserving blockchains (Monero and Zcash). We identified these blockchains’ addresses in 40 channels, and through manual verification, we confirmed that 29 of them are used for donations. Analyzing the BTC wallets, we find they received 115 transactions, totaling 0.5 BTC (\$$\approx\$13,000$). Performing the same analysis on Ethereum, we find that the wallets received 42 transactions, amounting to 5.5 ETH (\$$\approx\$9,000$).
Custom websites. Analyzing the URLs and examining messages within the conspiracy channels, we observed frequent promotion of custom e-commerce sites or personal websites. Inspecting these websites, we discover that many of them feature dedicated donation sections with blockchain addresses or various payment options. Thus, we perform a raw analysis to estimate the magnitude of the phenomenon. In particular, we look for URLs containing the words: shop, products, store, produkt, collections, donate, donations or support as third-level domain or that have these words in the URL’s path. As a result, we find 152,680 URLs (24,980 unique) matching our definition. Unfortunately, we can not validate or analyze this huge amount of websites since it requires a heavy manual effort or to build custom parsers.
@@ 6 Discussion
False positive channels. Validating channels present in the detected communities Sec. 4.2.4, we classify 121 out of 414 channels as false positives. In particular, we identified 28 channels associated with the American far-right or supporting the former Brazilian president Bolsonaro. Although the messages on these channels do not explicitly endorse or promote conspiracy theories, they often share a similar underlying narrative. Additionally, the relationship between conspiracy theories and the American far-right [88] or Bolsonaro [73] is well established. Furthermore, 24 channels cover topics related to religion and self-improvement. According to social and psychological scientific works [42, 24, 22], there exists a connection between these topics and conspiracy theories. Indeed, conspiracy theories often provide a sense of belonging to a community and the perception of control over complex and seemingly chaotic events. Moreover, some conspiracy theories, like QAnon, incorporate seemingly religious elements such as beliefs in a secretive satanic elite, apocalyptic prophecies, and the involvement of divine or demonic forces. Surprisingly, 15 channels are linked to cryptocurrency. This seemingly unusual connection is backed by a recent study, which reveals that cryptocurrency holders often exhibit strong anti-establishment sentiments and a pronounced tendency to believe in conspiracy theories [67, 68]. Lastly, there are 13 channels dedicated to alternative news. These channels provide unconventional viewpoints that often challenge traditional narratives.
@@
Detecting new monetization techniques. In this work, we focus on three main monetization methods used by conspiracy channels and mention other methods we’ve observed. However, new techniques may appear in the future. Based on our experience, we suggest the following ways to identify potential new techniques. Using the monetization messages collected in this study, an LLM can be fine-tuned to detect these messages more effectively than the standard ChatGPT we used. By monitoring the links these channels share and analyzing the most frequently shared ones with the fine-tuned model, it will be possible to identify new platforms used for monetization. Finally, many monetization strategies used by conspiracy channels are similar to those used by web content creators. Therefore, if web content creators start using new strategies, it is worth checking if conspiracy channels are also adopting them.
@@
The risks of conspiracy narratives. Our analysis reveals a clear distinction between the products promoted by conspiracy theory channels on Amazon compared to other online marketplaces. Indeed, while Amazon features standard products like books, masks, and water filters, we find a range of questionable products (e.g., 5G shields, EMF stone protectors, and healing wands) on eBay, Etsy, and Teespring. The distinction is likely attributed to the different content policies of these platforms. Indeed, Amazon upholds a more rigorous content policy compared to the other services. Nevertheless, it is important to emphasize that these products are not inherently harmful. Indeed, the concern lies in the narratives and promotions associated with these items. For example, ordinary substances like Sodium Chloride and Chlorine Dioxide, commonly found in pharmacies, are promoted in four channels as essential components to prepare the so-called "Miracle Mineral Supplement", which is reported as a miraculous cure for various diseases, including cancer and HIV. Similarly, we discover 35 channels sharing links to buy seemingly innocuous white pine needles at an exorbitant price of \$150 on Etsy. The concern lies in the accompanying message, which presents a guide for COVID-19 survival. This guide discourages seeking medical care in hospitals and suggests homemade remedies, including tea prepared with the costly pine needles mentioned above.
@@
Marketplace policies. Although Amazon has a rigorous policy about the items sold and actively operates to ban QAnon merchandising [77], it has a less strict policy about book content. Quoting Amazon policies: As a bookseller, we believe that providing access to the written word is important, including content that may be considered objectionable. [11]. The combination of this less strict policy and the simplicity of self-publishing on Amazon allowed conspiracy theorists to spread their ideas and monetize through book sales. Indeed, nearly 50% of the links point to books and almost half of them include an affiliate tag. Looking at the most frequently occurring authors of books promoted using non-affiliate links, we discover three relatively obscure German writers: VEIBZ (3,091 occurrences), MERKSAM (1,492 occurrences), and EBURD (813 occurrences). Their books explore several conspiracies with the goal of revealing the truth on subjects such as: "NASA & Elon Musk – They lie & cover-up," "CERN & its satanic roots," "HAARP & CERN use Alien Tech." Fig. 5 shows the cover of the three most shared books.
@@
Concerning the promoted items, there are also problems related to the transparency of the activity and compliance with the referral programs. Indeed, according to the partnership agreement of Amazon and Ebay [10] [43], a partner—a participant that can generate referral links and earn commissions—has to clearly disclose their partnership. Moreover, near each affiliated link should be a disclosure such as "(paid link)," "#ad," or "#CommissionsEarned". This information is needed to inform the customers that a conflict of interest exists on the promoted item.

### v3

@@ 4.3.1 Longitudinal analysis
An interesting aspect to explore is how conspiracy communities have evolved over time. We analyze this dimension in Fig. 2, which shows the number of channels created daily on Telegram. The Figure is divided into five parts: the first four charts ($a$,$b$,$c$,$d$) analyze each community, while the last one ($e$) compares the four conspiracy communities aggregated against the other Telegram communities. The chart shows that the creation of conspiracy channels is not evenly distributed over time. Instead, we find two spikes in channel creation. The first one begins around mid-March 2020, reaches a peak in May, and starts declining until June 2020. Instead, the second spike is much steeper and goes from January 6, 2021 to the middle of March 2021.
@@ 5.1 Donation platforms
Total 123,339 2,819 1,175,557 136,710
The first monetization strategy we analyze consists of asking the subscribers for donations to support the channel and its activity. As a first step, we identify the most popular donation platforms that can be used for this purpose. We manually collect them by analyzing the results of Google queries such as: top donation platforms. Given the presence of language-specific communities, we conducted country-specific queries containing different languages and keywords. Moreover, to avoid search results on Google from being affected by our browsing history or geo-location, we conduct queries using a VPN. In this way, we collect 31 donation services. Then, we analyze the presence of donation platforms in the URLs shared in the Extended Conspiracy URLs Dataset.
@@
After filtering out messages classified by the LLM as discrediting the linked donation page, we identified 124,610 URLs from 15 services, shared across 5,585 channels, accounting for 31.3% of our dataset. The most used platforms by number of URLs are Paypal/donate [68] with 83,512 URLs (67% of the donation URLs), Patreon with 31,041 URLs (24.9% of the donation URLs), BuyMeACoffee with 3,157 (2.5%), SubScribeStar with 2,604 (2.1%), DonorBox with 1,550 (1.2%) and Ko-Fi with 1,475 (1.2%).
To estimate the potential earnings of the Conspiracy Channel Dataset on these platforms, we create a custom crawler using the Selenium Framework [74]. We crawl all the URLs from the 6 most used donation platforms in our dataset, covering 98.9% of the URLs [^1] We conducted the crawling of donation, crowdfunding, and affiliate URLs in September 2023.. Unfortunately, PayPal/donate and DonorBox do not provide information about donation amounts and the number of donors. Thus, by scraping the web pages of these services, we could only determine the number of unique profiles still reachable. Instead, we obtained information regarding the number of donors and donated amounts for most profiles on the other services. Specifically, for Patreon and SubscribeStar, we retrieve the number of donors and the total amount of money donated in the last month. As for BuyMeACoffee and Ko-Fi, we collect data on the number of donors and the total amount of money raised by the campaigns’ creators.
@@ 5.2 Crowdfunding and Fundraising services
Thus, we use the same approach as the previous subsection to discover these cases (Sec. 5.1), asking the GPT-4o model to determine the author’s intention behind each message (prompt in Appendix, Tab. 8). Out of 7,499 distinct messages containing a URL to a crowdfunding project, the model classifies 7,319 messages (97.6%) as promoting campaigns while 180 as messages against the linked campaign. Also, in this case, to validate the model, we randomly sampled 200 messages, 100 marked as positive and 100 as negative. Then, we manually label these messages as previously done in Sec. 5.1. Here, we identified 145 positive samples and 55 negative samples. Comparing the model results with the manually annotated labels, we observed the model tends to classify positive messages as negative, similar to the previous task. Specifically, the model incorrectly classified 45 messages endorsing crowdfunding campaigns as negative. After reviewing the misclassified messages, we find many authors initially discredit the GoFundMe platform, encouraging their audience not to use it, and then endorse a crowdfunding campaign on the GiveSendGo platform. This pattern was common among American far-right channels, which frequently complain about GoFundMe blocking their projects. Given that the model provides a conservative estimation and considering the relatively low number of misclassified messages (180), we used this model for our classification. Accordingly, we discarded 180 unique messages (357 including duplicates) linking to crowdfunding platforms.
As further validation, we manually examine the crowdfunding campaigns that raised more than \$200K, counting 98 different projects, to determine if they are related to conspiracy theories. Three researchers independently analyzed these projects, and a campaign was labeled as related to conspiracy theories only if there was unanimous agreement. Among the 107 campaigns, we identified 51 projects (47.66%) that, despite being supported by conspiracy channels, have goals unrelated to conspiracy theories.
@@
Tab. 4 reports the number of URLs, different projects, money raised, and number of backers for the top eight services by number of URLs. As it is possible to note, this strategy is the most remunerative, collectively funding conspiracy theorists with almost \$71M. Analyzing the content of the campaigns and find that they typically fall into these categories:
Campaigns supporting far-right. We identified several fundraising campaigns for far-right projects, mainly raising funds for legal costs related to the January 6th Capitol riot. Three of these campaigns alone  [50, 71, 49] collected \$188,215 from 2,920 backers. We find that many of these campaigns are hosted on Givesendgo, a platform well known for promoting extremist content [89].
Campaigns about COVID-19. We also find GiveSendGo campaigns frequently linked to COVID-19, including those for the Freedom Convoy 2022 [43], a protest movement against COVID-19 policies. Similar GoFundMe campaigns redirect to refund pages due to the platform’s ban on content promoting violence [42]. Looking at the news, we discover that one of these campaigns was the biggest gainer of the platform, raising \$9,748,168 from 112,943 donors [16]. Additionally, Indiegogo and Kickstarter host projects funding documentaries about COVID-19 [27, 78, 92]. These platforms have stricter content policies, but moderating pseudo-scientific claims remains challenging due to freedom of expression concerns.
Scam campaigns. We detect crowdfunding campaigns that are outright scams. An example is a Fundrazor campaign [58] falsely claiming to raise funds for starving children in Venezuela. It collected \$39,500 from 602 donors before being shut down, and a banner on the page states that Save The Children has confirmed they are not associated. Another scam [48] on Indiegogo raised over \$7.4K for a portable air cleaner claiming to purify the air with a high-frequency generator before being closed as the creator claimed that Silicon Valley internet companies had hindered the project.
@@ 5.3 Affiliate Programs
Conspiracy channels adopting this strategy use affiliate programs from e-commerce platforms to earn a commission. Thus, as a first step, we identify the most popular e-commerce platforms they can utilize for this purpose. To this end, we leverage SimilarWeb to retrieve all the 42 services reported in the global rank for the eCommerce & Shopping category. Then, we extract URLs containing the domain of these platforms from the Extended Conspiracy URLs Dataset, finding 70,149 URLs from 28 platforms. Among the top 4 platforms, covering 98.1% of e-commerce links, only Amazon and eBay offer an affiliate program. In this program, a participant, known as partner, can generate unique links pointing to products on the platform that embed his unique identification number. The partner earns a commission for any purchases made by users who land on the platform through his link. In particular, Amazon’s partners earn a commission, between 1% and 12% depending on the categories of the product and the location of the targeted market (e.g., amazon.com, amazon.de) from all the items bought in the next 24 hours from arrival through the link [8]. Similarly, on eBay, a partner earns a commission between 1% and 4% on the items purchased on the platform in the following 24 hours [37].
Overall, we find 63,445 URLs pointing to Amazon and 1,932 to eBay in the Conspiracy Channel Dataset. To detect URLs belonging to the affiliate program, we search for URLs containing the parameters tag= for Amazon and campid= for eBay. Surprisingly, we discover that 34,412 (54.2%) of Amazon’s URLs and 131 (6.8%) on eBay are affiliated links granting commissions to the channels and that 1,810 channels (10.2% of our dataset) leverage this strategy. Tab. 5 provides a summary of the metrics of e-commerce platforms.
@@ 5.4 Monetization Analysis
Monetization links promoted by multiple channels. The analysis of Sabmyk affiliate links highlights the sharing of the same referral link by multiple channels. To study this phenomenon in more depth, we conducted a broader analysis, including donations and crowdfunding projects. Fig. 3 shows the cumulative distribution function (CDF) depicting the fraction of referral IDs (orange), crowdfunding (blue), and donation projects (green) shared in channels. The main plot provides an overview of the entire distribution, while two subplots show two different levels of zoom. By examining the zoomed CDFs, we can observe that most referral links, donation projects, and crowdfunding projects are typically shared by only a few channels. Specifically, we find that 65.4% of affiliate IDs, 56.8% of donation projects, and 55.8% of crowdfunding projects are shared by only a single channel. The figure also highlights that affiliate links are shared by fewer channels than donation and crowdfunding projects, with almost 80% of referral links shared by at most two channels. This can be expected as, while affiliate links are primarily beneficial to the channel owner, crowdfunding and donation projects may sometimes support broader causes and may be shared by channel owners even if they are not the direct beneficiaries of the funds received. Looking at the tail of the overall distribution, we find other interesting insights. Examining crowdfunding projects, we find that the most widely shared is a highly popular project supporting the Freedom Convoy, promoted on GiveSendGo [91]. This project is shared by 247 conspiracy channels and has raised \$9.7 million. For donation projects, we identified a donation link that was shared by 844 channels. This link is associated with the Corona_Fakten channel, which focuses on disseminating news and controversial theories about COVID-19. The channel frequently includes the donation link in its posts, and because its content is widely shared and forwarded by many other channels, the donation link has spread across hundreds of channels. What is more surprising instead, is the presence of a group of over 549 channels promoting Amazon products using the same affiliate ID. Our analysis reveals this group of channels is entirely located in the German community and is actively involved in promoting various conspiracy theories, predominantly related to COVID-19.
@@
Shift in platform usage. Fig. 7 and Fig. 8, in the Appendix, show, respectively, the proportion of donation and crowdfunding platforms used by each community monthly. Examining donation platforms, we observe distinct preferences across different linguistic communities. The German community predominantly uses PayPal. In contrast, Patreon has been the most used platform in the Neo-Latin community over the last year. The English community initially used Patreon and PayPal most frequently, but in the past 18 months, there has been a significant diversification, with multiple platforms being utilized. For crowdfunding, PayPal Pools was the dominant platform among both the German and Neo-Latin communities. However, its popularity decreased until it was discontinued in September 2021 and replaced by GoFundMe. The English community displayed a different trend. GoFundMe was consistently the preferred platform, as PayPal Pools was never widely adopted. Around October 2020, GiveSendGo began to rise in popularity, accounting for 75% to 90% of all URLs shared by the English community monthly. This shift may be attributed to GoFundMe stricter policies, as GiveSendGo promotes itself as having no moderation. We find evidence of this transition in the messages of the English community.
@@ 5.5 Other Strategies
In addition to the monetization strategies analyzed previously, conspiracy channels can exploit other sources of revenue.
@@
Blockchain addresses. In our analysis, we discover conspiracy channels asking for cryptocurrency donations in their descriptions. We use regular expressions to extract wallet addresses of the most popular blockchains (Bitcoin and Ethereum), as well as the prominent privacy-preserving blockchains (Monero and Zcash). We identified these blockchains’ addresses in 40 channels, and through manual verification, we confirmed that 29 of them are used for donations. Analyzing the BTC wallets, we find they received 115 transactions, totaling 0.5 BTC ($\approx\$13,000$). Performing the same analysis on Ethereum, we find that the wallets received 42 transactions, amounting to 5.5 ETH ($\approx\$9,000$).
Custom websites. Analyzing the URLs and examining messages within the conspiracy channels, we observed frequent promotion of custom e-commerce sites or personal websites. Inspecting these websites, we discover that many of them feature dedicated donation sections with blockchain addresses or various payment options. Thus, we perform a raw analysis to estimate the magnitude of the phenomenon. In particular, we look for URLs containing the words: shop, products, store, produkt, collections, donate, donations or support as third-level domains or that have these words in the URL’s path. As a result, we find 152,680 URLs (24,980 unique) matching our definition. Unfortunately, we can not validate or analyze this huge amount of websites since it requires a heavy manual effort or to build custom parsers.
@@ 6 Discussion
False positive channels. Validating channels present in the detected communities Sec. 4.2.4, we classify 121 out of 414 channels as false positives. In particular, we identified 28 channels associated with the American far-right or supporting the former Brazilian president Bolsonaro. Although the messages on these channels do not explicitly endorse or promote conspiracy theories, they often share a similar underlying narrative. Additionally, the relationship between conspiracy theories and the American far-right [72] or Bolsonaro [35] is well established. Furthermore, 24 channels cover topics related to religion and self-improvement. According to social and psychological scientific works [36, 21], there exists a connection between these topics and conspiracy theories. Indeed, conspiracy theories often provide a sense of belonging to a community and the perception of control over complex and seemingly chaotic events. Moreover, some conspiracy theories, like QAnon, incorporate seemingly religious elements such as beliefs in a secretive satanic elite, apocalyptic prophecies, and the involvement of divine or demonic forces. Surprisingly, 15 channels are linked to cryptocurrency. This seemingly unusual connection is backed by a recent study, which reveals that cryptocurrency holders often exhibit strong anti-establishment sentiments and tend to believe in conspiracy theories [57]. Lastly, there are 13 channels dedicated to alternative news. These channels provide viewpoints that often challenge traditional narratives.
@@
Detecting new monetization techniques. In this work, we focus on three main monetization methods used by conspiracy channels and mention other methods we’ve observed. However, new techniques may appear in the future. Based on our experience, we suggest the following ways to identify potential new techniques. Using the monetization messages collected in this study, an LLM can be fine-tuned to detect these messages more effectively than the standard ChatGPT we used. By monitoring the links these channels share and analyzing the most frequently shared ones with the fine-tuned model, it will be possible to identify new platforms used for monetization. Finally, many monetization strategies used by conspiracy channels are similar to those used by web content creators. Therefore, if they adopt new strategies, it is worth checking if conspiracy channels start using them.
@@
The risks of conspiracy narratives. Our analysis reveals a clear distinction between the products promoted by conspiracy theory channels on Amazon and other online marketplaces. Indeed, while Amazon features standard products like books, masks, and water filters, we find a range of questionable products (e.g., 5G shields, EMF stone protectors, and healing wands) on eBay, Etsy, and Teespring. The distinction is likely attributed to the different content policies of these platforms. Indeed, Amazon upholds a more rigorous content policy compared to the other services. Nevertheless, it is important to emphasize that these products are not inherently harmful. Indeed, the concern lies in the narratives and promotions associated with these items. For example, ordinary substances like Sodium Chloride and Chlorine Dioxide, commonly found in pharmacies, are promoted in four channels as essential components to prepare the so-called "Miracle Mineral Supplement", which is reported as a miraculous cure for various diseases, including cancer and HIV. Similarly, we discover 35 channels sharing links to buy seemingly innocuous white pine needles at an exorbitant price of \$150 on Etsy. The concern lies in the accompanying message, which presents a guide for COVID-19 survival. This guide discourages seeking medical care in hospitals and suggests homemade remedies, including tea prepared with the costly pine needles mentioned above.
@@
Marketplace policies. Although Amazon has a rigorous policy about the items sold and actively operates to ban QAnon merchandising [12], it has a less strict policy about book content. Quoting Amazon policies: As a bookseller, we believe that providing access to the written word is important, including content that may be considered objectionable. [9]. The combination of this less strict policy and the simplicity of self-publishing on Amazon allowed conspiracy theorists to spread their ideas and monetize through book sales. Indeed, nearly 50% of the links point to books, and almost half of them include an affiliate tag. Looking at the most frequently occurring authors of books promoted using non-affiliate links, we discover three relatively obscure German writers: VEIBZ (3,091 occurrences), MERKSAM (1,492 occurrences), and EBURD (813 occurrences). Their books explore several conspiracies with the goal of revealing the truth on subjects such as: "NASA & Elon Musk – They lie & cover-up," "CERN & its satanic roots," "HAARP & CERN use Alien Tech." Fig. 5 shows the cover of the three most shared books.
@@
Concerning the promoted items, there are also problems related to the transparency of the activity and compliance with the referral programs. Indeed, according to the partnership agreement of Amazon and Ebay [8] [37], a partner—a participant that can generate referral links and earn commissions—has to clearly disclose their partnership. Moreover, near each affiliated link should be a disclosure such as "(paid link)," "#ad," or "#CommissionsEarned". This informs customers that there is a conflict of interest in the promoted item.